### v2

# I Learn Better If You Speak My Language: Understanding the Superior Performance of Fine-Tuning Large Language Models with LLM-Generated Responses

Abstract
This paper explores an intriguing observation: fine-tuning a large language model (LLM) with responses generated by a LLM often yields better results than using responses generated by humans. We conduct an in-depth investigation to understand why this occurs. Contrary to the common belief that these instances is simply due to the more detailed nature of LLM-generated content, our study identifies another contributing factor: an LLM is inherently more "familiar" with LLM generated responses. This familiarity is evidenced by lower perplexity before fine-tuning. We design a series of experiments to understand the impact of the "familiarity" and our conclusion reveals that this "familiarity" significantly impacts learning performance. Training with LLM-generated responses not only enhances performance but also helps maintain the model’s capabilities in other tasks after fine-tuning on a specific task.
I Learn Better If You Speak My Language: Understanding the Superior Performance of Fine-Tuning Large Language Models with LLM-Generated Responses
Xuan Ren xuan.ren@adelaide.edu.au                        Biao Wu biaowu165534@gmail.com                        Lingqiao Liuthanks: Corresponding author. lingqiao.liu@adelaide.edu.au
1 Introduction
Recent research has demonstrated that a large language model (LLM) can generate training data for another LLM Dai et al. (2023); Edwards et al. (2022); Møller et al. (2024); Guo et al. (2023); Ubani et al. (2023); Agrawal et al. (2023); Xu et al. (2023); Kieser et al. (2023); Taori et al. (2023); Peng et al. (2023); Xu et al. (2024). This approach offers a method for transferring knowledge from a larger model to a smaller one, or for creating supplementary training materials, such as rationales Zhang et al. (2024); Kang et al. (2023); Li et al. (2022), or refined step-by-step reasoning Hsieh et al. (2023); Ho et al. (2022); Magister et al. (2023); Fu et al. (2023).
The quality of data generated by large language models (LLMs) is considered to be very high and conducive to effective trainingLee et al. (2023). The prevailing view often attributes the successful training results of LLM-generated data to the fact that LLMs provide more details than human-annotated data, such as ’chain-of-thought-style’ reasoning processHsieh et al. (2023); Ho et al. (2022); Magister et al. (2023); Fu et al. (2023). However, our experiments have found numerous counterexamples. As shown in Figure 1, target labels that include more details and utilize chain of thought styles do not necessarily result in better training outcomes. Through experiments on various tasks, we have observed that when LLMs directly generate target labels, the training results are often superior to those obtained by transforming ground truth into more detailed, step-by-step training. This indicates that whether the answer generation is executed in the LLM’s own style or is based on a rewritten human style is a key factor affecting training outcomes.
Figure 1: This figure illustrates the training outcomes associated with various data generation methods. As demonstrated by the examples shown, the inclusion of additional details does not necessarily translate into enhanced performance.
This observation motivates us to investigate the underlying reason for this phenomenon. We found that whether the LLM is familiar with the target label is a crucial factor influencing model training.
Our main hypothesis are as follows:
  • Given the same question, an LLM is inherently familiar with responses generated by itself or other LLMs.
  • an LLM can better adapt to and leverage responses in its familiar styles, leading to improved outcomes.
The following section of this paper will examine the hypothesis through a series of experiments. We will measure ’familiarity,’ conduct ablation studies by creating different variants of LLM-generated responses, and evaluate their performance to understand various factors in the process. This will provide insights into our hypothesis.
2 Empirical Study Protocols
The purpose of this paper is to understand and investigate the aforementioned phenomenon rather than to propose a new method. Our study’s methodology is based on empirical analysis. We begin by introducing our study protocols, including the datasets and evaluation approaches used.
2.1 Evaluation approaches
Without loss of generality, we denote the training data as question and response pairs $(q,r)$. We will use a subset of the LLaMA2 benchmark Touvron et al. (2023) to create the training dataset. We will use a subset of the LLaMA2 benchmark [34] that contains human-labeled explicit reasoning processes to create the training dataset. The question and response pairs in these datasets are human-constructed. Additionally, we will use an LLM to generate new responses, following $r^{\prime}=LLM(q,P)$, where $LLM$ denotes the language model and $P$ denotes the prompt used to create the response. We will explore different $P$ in the following experiment. This creates a synthetic version of dataset $\{(q,r^{\prime})\}$.
Our training approach for the LLaMA2-13b-chatTouvron et al. (2023) and Mistral-7b-Instruct-v2Jiang et al. (2023) models involves specific adaptations to optimize performance. Notably, the warm-up period is set to 10% of the total training steps consistently across our experiments. For the LLaMA-13b-chat model training on the HumanEval dataset, we set the learning rate to 2e-4. This adjustment was necessary as lower rates proved insufficient for the model to effectively learn from this dataset. Given the small size of the training datasets, such as MBPP and HumanEval, we reduced the batch size to 10 to accommodate the limited data. In contrast, for general fine-tuning of the LLaMA2-13b-chat and Mistral-7b-Instruct-v2 models, we employ a default learning rate of 2e-5, a batch size of 32, and a warm-up period comprising 10% of the total steps. Additionally, we use a cosine learning rate schedule and fine-tune only the Q and V matrices of the LORAHu et al. (2021) parameters with a rank of 8. All models are trained and evaluated using half-precision.
We evaluate the fine-tuned model in two different ways. One is called in-domain evaluation, which assesses performance on the tasks of the training set. The other is called cross-task evaluation, which evaluates performance across tasks on other datasets. The purpose of using cross-domain evaluation is to test if the model can benefit from the training process of a similar task and if the model will forget its capability on other tasks after training on a specific task.
3 Experiment setup
3.1 Dataset and Evaluation
We utilized the LLaMA2 benchmark Touvron et al. (2023) to select five tasks that contain human-labeled explicit reasoning. These were evaluated based on accuracy to assess the effectiveness of different data construction methods. The term ’dataset with explicit reasoning’ refers to datasets that include a detailed reasoning process in their annotations. This category includes GSM8KCobbe et al. (2021), MATHHendrycks et al. (2021), HumanEvalChen et al. (2021), MBPPAustin et al. (2021), and ECQA Aggarwal et al. (2021), where the ground truth comprises a complete human-labeled problem-solving process or rationale, not just the final answer. For example, the ground truth in GSM8K involves comprehensive mathematical reasoning, and the coding datasets’ ground truths consist of complete codes, which are logical sequences by nature. We included coding datasets like MBPP and HumanEval in this category, as developing code inherently involves logical reasoning, even though it may not always be presented in the form of a traditional chain of thought. Our model was trained on one dataset and subjected to in-domain testing on the same dataset, as well as cross-domain testing on other datasets. Our selection includes all coding and mathematics datasets within the LLaMA2 benchmark. We chose to use the ECQA dataset because, unlike the CQATalmor et al. (2018) dataset included in the LLaMA2 benchmark, ECQA includes rationales provided by human annotators.
Code: We evaluated coding ability using the HumanEval Chen et al. (2021) and MBPP Austin et al. (2021) datasets. We report the pass@1 accuracy for each dataset. Note that the HumanEval dataset contains only 164 testing data points with no training data. We divided the training task into two parts: in the first part, we trained on the first 82 data points and tested on the last 82. In the second part, we reversed this order. Ultimately, we averaged the test results from both parts. For the MBPP-full dataset, as opposed to the MBPP-sanitized version which includes fewer training data points, we utilized all available training data (374 data points) and testing data (500 data points)
Commonsense Reasoning: For Commonsense QA task, we selected the ECQA Aggarwal et al. (2021) dataset, which provides explanations for the answers from CQATalmor et al. (2018) dataset. We combined these explanations with the multiple-choice answers to form the ground truth training data. We used the first 1000 training and test data points for training and evaluation.
Math: We conducted training and evaluations on the GSM8K Cobbe et al. (2021) and MATH Hendrycks et al. (2021) datasets. The MATH dataset encompasses several categories. We conducted training and evaluations exclusively within the algebra category. For the evaluation, we developed a script that compares the ground-truth answer with the model’s final prediction. This script accurately evaluates responses that are numerical. However, in the MATH dataset, some answers contain inequalities or complex mathematical expressions. We filtered out these data points and used only those with numerical answers for both training and evaluation. After excluding data points with non-numerical answers, we used the first 1000 training data points for training and all the testing data points for testing. As a result, we obtained 1000 training and 1314 testing data points for GSM8K, and 1000 training and 752 testing data points for MATH (algebra).
Evaluation Script:
To facilitate automatic evaluation across all models and datasets, we developed a specialized prompt format requiring models to conclude their predictions with ’Final Answer: answer.’ For instance, the output should read ’Final Answer: 100’ for an answer of 100. Our evaluation script extracts the text following the ’Final Answer:’ substring. For the ECQA task, it verifies if the first text after ’Final Answer:’ correctly matches one of the answer options. For mathematical problems including the GSM8K and MATH datasets, the script checks if the number after the substring ’Final Answer:’ aligns with the ground truth number. We utilize the built-in testing script for the HumanEval dataset to ensure result accuracy and employ our custom scripts for all other tasks. By adapting the training data’s question and answer formats to this style, our models maintain consistent response formatting across different tasks. The evaluation scripts for HumanEval Chen et al. (2021) and MBPP Austin et al. (2021) require answers in a specific format. Models that have not been trained on this format struggle to follow it correctly, even when provided with in-context examples. Consequently, we do not conduct cross-domain evaluations on the MBPP and HumanEval datasets.
4 Observations and Hypothesis:
4.1 Observation 1: Superior Performance of GPT-4 Over Human-Annotated Labels
Method Training Dataset and Model Type GSM8K Math Algebra ECQA HumanEval MBPP
Groundtruth GSM8K, Mistral 0.434 0.162 0.594
GPT-4 Answer Directly 0.597 0.246 0.597
Groundtruth GSM8K, Llama 0.364 0.141 0.565
GPT-4 Answer Directly 0.428 0.128 0.575
Groundtruth Math algebra, Mistral 0.264 0.206 0.554
GPT-4 Answer Directly 0.553 0.302 0.608
Groundtruth Math algebra, Llama 0.36 0.126 0.575
GPT-4 Answer Directly 0.35 0.150 0.561
Groundtruth ECQA, Mistral 0.258 0.134 0.68
GPT-4 Answer Directly 0.462 0.223 0.722
Groundtruth ECQA, Llama 0.132 0.0798 0.631
GPT-4 Answer Directly 0.379 0.156 0.656
Groundtruth HumanEval, Mistral 0.3625 0.191 0.583 0.323
GPT-4 Answer Directly 0.313 0.163 0.581 0.366
Groundtruth HumanEval, Llama 0.0705 0.083 0.528 0.146
GPT-4 Answer Directly 0.1245 0.105 0.552 0.159
Groundtruth MBPP, Mistral 0.392 0.176 0.527 0.276
GPT-4 Answer Directly 0.399 0.186 0.568 0.354
Groundtruth MBPP, Llama 0.328 0.132 0.556 0.2
GPT-4 Answer Directly 0.351 0.140 0.574 0.202
Table 1: Comparison of human-annotated data versus data generated directly by GPT-4. We employ GPT-4 to generate responses using only the questions from the training datasets. For the ECQA task, we provide GPT-4 with gold labels (excluding rationales) to guide its response generation. In-domain performance is highlighted in grey. Data points significantly lower in accuracy by 15% or more relative to another method are marked in red.
Our study is inspired by a compelling observation: when a sophisticated LLM, such as GPT-4, is used to generate responses to questions, and these questions along with their generated responses are used as a synthetic dataset, fine-tuning smaller LLMs (we use both Llama2-13B-Chat and Mistral-7B-Instruct-v2) on this synthetic dataset often yields better performance than using the original dataset with human-provided answers. The detailed experimental results are presented in Table 1. As demonstrated, the majority of the low-performance data points (14 out of 15) marked in red reflect accuracies when training with human-annotated ground truth. Using synthetic datasets consistently yields higher in-domain performance compared to human-answered datasets. Notably, we observe a significant improvement in math-related tasks, often exceeding a 10% absolute increase in in-domain performance. Additionally, the model trained on the synthetic data achieves higher cross-domain performance. This observation motivates us to investigate the underlying reasons.
4.2 Observation 2: significantly lower perplexity of GPT-4 and other LLMs Over Human-Annotated Labels
Figure 2: Average Perplexity Comparison
We measure the perplexity of responses as assessed by the target (smaller) LLMs. An interesting observation, shown in Figure 2, is that, given the input questions from the responses generated by GPT-4 attain significantly lower perplexity compared to those provided by humans. To explore whether this phenomenon is unique to GPT-4 or common among LLM responses, we employed a zero-shot approach with the Mistral-7B-Instruct-v2 model and the LLaMA-13B-chat model. These models answered the same questions from the training set. As shown in Figure 2, their perplexity was also significantly lower than that of the human-labeled dataset. Lower perplexity indicates that LLMs are more familiar with responses generated by themselves or other LLMs than with human-generated responses.
4.3 Hypotheses
From Observation 2, we notice that given the same question, the average perplexity of responses generated by LLMs is consistently lower than the perplexity of ground truth annotated by humans. This observation leads us to pose our first hypothesis: When generating target labels for the same question, one LLM inherently more familiar with data generated by another LLM. We not only observed consistently lower perplexity from GPT-4 across tasks in Observation 2 but also noted that LLMs trained on GPT-4 generated responses consistently outperform the same LLMs trained using human-annotated data. This phenomenon leads us to propose our second hypothesis: An LLM performs better when trained on data it is familiar with.
5 Investigation of the proposed hypotheses
Method Training Dataset and Model Type GSM8K Math Algebra ECQA Perplexity token length
Groundtruth GSM8K, Mistral 0.434 0.162 0.594 4.54 128.88
GPT-4 Answer Directly 0.597 0.246 0.597 2.35 179.94
Step by step 0.574 0.224 0.591 2.27 282.9
Step by step transformation of GT 0.499 0.190 0.6 2.67 235.892
Groundtruth GSM8K, Llama 0.364 0.141 0.565 3.27 128.88
GPT-4 Answer Directly 0.428 0.128 0.575 2.30 179.941
Step by step 0.402 0.134 0.553 2.19 290.92
Step by step transformation of GT 0.394 0.146 0.573 2.43 242.22
Groundtruth Math algebra, Mistral 0.264 0.206 0.554 5.83 163.36
GPT-4 Answer Directly 0.553 0.302 0.608 2.20 328.76
Step by step 0.533 0.239 0.583 2.249 451.99
Step by step transformation of GT 0.444 0.210 0.577 2.65 370.35
Groundtruth Math algebra, Llama 0.36 0.126 0.575 5.15 163.36
GPT-4 Answer Directly 0.35 0.150 0.561 2.59 328.76
Step by step 0.344 0.145 0.593 2.55 450.08
Step by step transformation of GT 0.376 0.141 0.565 2.81 369.06
Groundtruth ECQA, Mistral 0.258 0.134 0.68 51.3 73.51
GPT-4 Answer Directly 0.462 0.223 0.722 6.35 176.54
Step by step 0.481 0.203 0.71 4.66 325.15
Step by step transformation of GT 0.462 0.194 0.681 6.89 218.748
Groundtruth ECQA, Llama 0.132 0.0798 0.631 11.96 73.51
GPT-4 Answer Directly 0.379 0.156 0.656 5.33 176.54
Step by step 0.363 0.140 0.648 4.46 337.30
Step by step transformation of GT 0.312 0.108 0.628 5.55 226.99
Table 2: Performance comparison of models trained on data constructed by different methods across various tasks. Data points are labeled as low performance when their accuracy is more than 15% lower relative to the highest accuracy achieved in the same dataset using the same model.
In this section, we conduct an in-depth investigation about the aforementioned hypotheses. Our investigations are organized into two parts. In Section 5.1, we examine to what extend the familiarity contribute to higher performance of using LLM generated responses. Particularly, we conduct experiments to ablate the factor that LLM could introduce additional reasoning details to the original human-provided response. In Section 5.2, we experiment to determine if a model, when used alone without the assistance of a more advanced model, can rewrite ground truth into a form it is more familiar with, thereby enhancing training outcomes. The advantage of this experimental design is that it allows us to evaluate the effect of familiarity while removing any benefits that might arise from using a more advanced language model.
5.1 Understand the impact of “familiarity”
Experiment 1: Is observation 1 only rooted from more detailed reasoning steps provided by the LLM other than familiarity?
Our observation 1 suggests that using GPT-4 generated data leads to better training effectiveness than using human-annotated ground truth. Our second hypothesis posits that this enhanced learning by LLMs on LLMs responses could be attributed to familiarity with the data. It is a common belief that GPT-4 often includes more detailed content than human annotators, which could contribute to improved training outcomes. However, we question whether the superior results from GPT-4 are solely due to the addition of more details, or if familiarity plays a significant role as well.
To explore whether familiarity is a crucial factor, we designed the following experiment. We employed different methods to have GPT-4 generate data in a detailed, step-by-step manner. For each dataset, we established a control group: one group involved GPT-4 transforming ground truth into its own detailed style, while the other had GPT-4 mimic the style of the human-labeled ground truth, adding detail to build upon it in a step-by-step style. As shown in Table 2, in the vast majority of cases, the training effectiveness of the first group, which used GPT-4’s own style, was superior to that of the second group, which followed the human-annotated style. In addition, it is noteworthy that using the Mistral 7B model for training on math algebra, converting ground truth into step-by-step answers was even detrimental. Both of the step by step groundtruth transformation methods achieves accuracy below 24% on Math Algebra, much lower than 30.2% achieved by ’GPT-4 answer directly’(Gpt-4 answer directly represents employing GPT-4 to generate responses using only the questions from the training datasets. For the ECQA task, we provide GPT-4 with gold labels (excluding rationales) to guide its response generation.) We speculate that one possible reason for this is that adding ’more than enough details’ may sometimes complicate the model’s reasoning process. We have included additional experiments involving various GPT-4 generated variants in the appendixA, which generally conform to the patterns described above. For clarity and conciseness, Table 2 only includes 2 examples of the step-by-step variants.
After preliminary validation of our hypotheses, we made another intriguing observation: According to Table 2, the ’GPT-4 answer directly’ dataset, characterized by a shorter average token length of target labels compared to both ’step by step’ and ’step by step transformation of GT’ datasets, suggests fewer details in its content. Despite this, the performance of the data directly generated by GPT-4 frequently ranks among the best of all GPT-4 answer variants. Meanwhile, other data generation methods exhibited significantly lower performance than GPT-4’s direct answers on at least two tasks, as highlighted in red in Table 2. This suggests that, given a problem, the data directly generated by GPT-4 might yield the good training results. Artificially defined prompts that add information to the ground truth may lead to the generation of target labels that do not conform to the LLM’s linguistic or logical preferences.
Experiment 2: Direct measuring the impact of familiarity
Method Model Type GSM8K GSM8K Perplexity Math Algebra Math Algebra Perpleity ECQA ECQA Perplexity
GPT-4 answer with lower perplexity Mistral 0.597 2.35 0.301 2.20 0.722 6.35
GPT-4 answer with higher perplexity 0.547 5.58 0.231 4.66 0.702 11.13
GPT-4 answer with lower perplexity LLaMA 0.428 2.30 0.150 2.59 0.656 5.33
GPT-4 answer with higher perplexity 0.38 4.19 0.131 4.35 0.654 7.69
Table 3: Comparison between the GPT-4 answer with higher perplexity Vs. lower perplexity
This section introduces a method for assessing the influence of style ’familiarity’. We direct GPT-4 to produce several varient of responses to the same query, grouping them into two distinct datasets based on their perplexity levels—high and low. We then evaluate and compare the learning outcomes using these two sets. One concern of this experiment is that GPT-4 may produce answers with different level of details. To eliminate this factor, we first let GPT-4 to generate one answer and then explicitly direct it to paraphrase the generated answer. In this way, we can create responses that vary in perplexity but maintain the same semantic meaning.
After training on those two sets of training data, we examine their performance, shown in Table 3. As seen, training on the higher-perplexity training set consistently shows slightly lower performance than the lower perplexity counterpart, particularly noticeable in the in-domain performance on two math datasets. For example, in the algebra dataset from Mistral, the lower perplexity set achieves a 30% accuracy rate, while the higher perplexity set only reached 23%. Since they are only differed in the style of the text, this result demonstrated that the familiarity of LLMs with the stylistic aspects of target labels significantly affects training outcomes.
From these findings, we conclude that LLMs perform poorly when trained on unfamiliar data, explaining the suboptimal training results with human-annotated data.
Discussion: Why are LLMs often familiar with responses generated from other LLMs? The familiarity observed when training on data generated by GPT-4 might stems from the overlap in training data used during the pre-training stages of these models. Many LLMs, including GPT-4, share similar pre-training data, developing an understanding of specific output styles and content.
5.2 Achieving high familiarity by exploring the target LLM generated responses
Math Commensense Reasoning NLI
Method Model Type GSM8K Math Algebra ECQA ESNLI
Groundtruth Mistral 0.434 0.206 0.68 0.814
Correct predicted sample + groundtruth 0.491 0.227 0.578 0.557
correct preidcted sample only 0.449 0.257 0.586 0.698
Style Transfered Groundtruth 0.51 19.6 0.696 0.845
Table 4: Performance Comparison when not using GPT-4. For the style-transferred ground truth dataset, only 73.7% and 24% of the training data are used for training, as explained in Section5.2
In the experiments described above, we validated two hypotheses. However, all the studies mentioned use GPT-4 to generate responses, chosen for its advanced capabilities to directly produce correct answers and reliably follow diverse instructions. However, there is a concern regarding the "familiarity" factor: would this argument still hold without the involvement of an advanced LLM? If valid, the argument should apply to some extent even without using advanced LLMs.
This section seeks to develop a method that enhances the “familiarity” of responses without requiring more advanced large language models (LLMs) than the target LLM.
We included ESNLICamburu et al. (2018) as an additional dataset to better compare different data construction methods. The e-SNLI dataset enhances the SNLIBowman et al. (2015) dataset by adding human-annotated explanations to each instance. These annotations offer insights into the reasoning behind the labels of entailment, contradiction, or neutral, improving model interpretability and aiding in the development of more robust inference systems.
Use the same model for train and groundtruth style transfer We devised a method to rewrite the ground truth data in a style that resembles language model-generated responses using the Mistral-7B-Instruct-V2 model. The experiment results are shown in Table 4. Initially, we posed questions from the training dataset and recorded the answers that were correctly predicted by Mistral. We documented two such correct predictions. These two correct responses were then used as in-context examples to guide Mistral on how to rewrite ground truth data. Specifically, we used the two recorded predictions as example target labels and their corresponding questions as example questions, while the actual ground truth served as the example ground truth. The in-context prompt informed Mistral that its task was to rewrite the ground truth in its own styles.
Subsequently, we presented all questions and their corresponding ground truths from the training dataset to Mistral for style rewriting. We recorded Mistral’s predictions and paired these with the original inputs to form a new training dataset, which we named "style transferred ground truth."
We noticed that even when provided with the ground truth, the Mistral model struggles to successfully rewrite answers for mathematical tasks. Therefore, we ran the rewritten math answers through our evaluation script and found that only 73.7% of the GSM8K answers and 24% of the math algebra answers passed the script. For math tasks, we only used the answers that passed the script for training. In the ECQA and ESNLI tasks, we allowed the Mistral model to rewrite the ground truth once and directly used all the transformed ground truth answers for training.
This experiment was conducted solely with Mistral and not with the Llama model, due to our observations that Llama2-13b-chat lacks sufficient in-context following capabilities required for successful ground truth style transformation. We show a failure example in AppendixD. As a comparative experiment, we created the ’correct predicted samples only’ dataset. For each sample in this dataset, we had the Mistral model perform zero-shot generation 5 times, and using an evaluation filter, we randomly selected one correct prediction that passed the filter to be paired with the original input as the target label. If the model failed to generate a correct answer in all 5 attempts, we did not include that data in the training dataset. The construction method for the ’correct predicted samples’ and ’correct predicted samples only’ datasets is similar. The only difference is that if the sample fails to produce a correct answer, we include the ground truth in the dataset as the correct answer.
As shown in Table 4, the Mistral model trained with style-transferred ground truth data outperformed the model trained with original ground truth data and other methods in all cases except for the Math Algebra dataset, where only 24% of the training data was used. To fairly compare the training effectiveness of the ’style transferred ground truth’ and the original ground truth, we recorded the IDs of the data used by the ’style transferred ground truth’. We identified a corresponding set of 240 original ground truth data points for comparison. We found that training with only these 240 data points resulted in a 17.4% accuracy using the original ground truth, which is lower than the 19.6% achieved by the ’style transferred ground truth’. This demonstrates that data processed using the ’ground truth style transfer’ method yields better training outcomes than data labeled by humans.
Interestingly, on the ECQA and ESNLI datasets, models trained with correct answers generated by Mistral in a zero-shot fashion performed worse than those trained directly with ground truth. By reading the ’correct predicted samples only’ dataset, we found that many instances classified as correct by the evaluation filter did not result from sound reasoning by the model. Although the final answers were technically correct, the underlying rationales were incorrect. Therefore, the effectiveness of the self-training method depends significantly on whether the evaluation filter used to select correct predictions can truly identify accurate responses.
6 A hybrid approach: creating responses using “minimal change” principal
Generally speaking, our experiments highlight two essential characteristics of effective responses. First, the target LLM must be "familiar with" the response. Second, the response must be accurate. The approach detailed in Section 5.2 utilizes outputs generated by the target LLM to ensure "familiarity" and employs "rewriting ground-truth response" to maintain accuracy. Familiarity is effectively assured as the response is produced by the model itself, often resulting in low perplexity. However, as shown in the experiments in Section 5.2, the strategy of "rewriting ground-truth response" frequently fails to guarantee correctness. In this section, we explore a hybrid approach where we use responses from the target LLM to achieve high familiarity and employ GPT-4 to make minimal adjustments to correct any inaccuracies, thereby ensuring both familiarity and correctness.
We call this approach “minimal change” and its construction process is as follows: For a given question, we first request a response from the target LLM, such as Llama2 or Mistral. We then instruct GPT-4 to copy the correct portions of the prediction and revise any erroneous parts of the response. When correcting errors, GPT-4 is required to make minimal changes to accurately adjust the answer  [^1] Please note that achieving such a process requires an advanced LLM, as the LLM needs to fully understand the solution and follows the “minimal change” principal. Note that this construction process ensures that the target LLM is “familiar” with the response as the response is largely generated from itself. The “minimal change” principal, on the other hand, minimizes the influence from the GPT-4.
This method involves minimal intervention by GPT-4 in the initial predictions to ensure that the original language style is largely retained. To effectively guide GPT-4 in producing these minimally changed training datasets, we used a specific prompt, outlined in the ’Minimum Change Prompt and Method’ section of the Appendix C, which includes two to three examples of minimal changes along with explanations for each modification.
Method Model Type GSM8K Math Algebra ECQA MBPP HumanEval Avg Perplexity
Groundtruth Mistral 0.434 0.206 0.68 0.276 0.323 15.03
GPT-4 Answer Directly 0.597 0.302 0.722 0.354 0.365 3.81
Minimum Change on Mistral Predictions 0.562 0.314 0.699 0.354 0.409 2.47
Minimum Change on LLaMA predictions 0.547 0.297 0.709 0.364 0.427 3.51
Groundtruth Llama 0.364 0.126 0.631 0.2 0.146 6.18
GPT-4 Answer Directly 0.428 0.150 0.656 0.2 0.158 3.58
Minimum Change on LLaMA predictions 0.433 0.166 0.649 0.2 0.213 2.75
Minimum Change on Mistral Predictions 0.402 0.161 0.647 0.218 0.183 3.32
Table 5: Comparing the experimental results of GPT4, groundtruth and minimum change
Results from our ’Minimum Change’ experiments are presented in Table 3. The data shows consistently lower perplexity compared to ground truth data across all datasets, indicating a high level of model familiarity with LLM-generated data. Furthermore, the performance across various setups was not only similar (with in-domain performance differences within 5% on most datasets) but also comparable to models trained with direct GPT-4 responses. Although the data was minimally modified based solely on the outputs from Mistral/LLaMA, theoretically, the logic and quality of the modified answers, as well as the level of detail, should be inferior to answers directly generated by GPT-4. Surprisingly, on the HumanEval and math algebra tasks, the performance of models trained with minimally changed data even surpassed those trained with data generated by GPT-4. On other datasets, the performance of the minimum change data was comparable to that of GPT-4. This underscores the effectiveness of training LLMs on data distributions familiar to them, demonstrating that minimal changes to maintain the original structure can be as effective as more detailed alterations. For more information about this experiment, please refer to the Appendix. B
7 Related Works
There are many studies related to using LLMs to generate training data for training models.
Self-Training: Several works utilize a model’s own predictions for self-training. For instance, STAR Zelikman et al. (2022) and REST Gulcehre et al. (2023) generates a dataset through sample production from the LLMs, subsequently utilizing these samples to enhance the LLMs via training. RESTem Singh et al. (2023) enhances model performance by using initial predictions, filtering for accuracy, and retraining the model with correct predictions. This iterative process improves the model’s accuracy over multiple cycles.
Data Augmentation: Numerous studies have investigated the use of synthetic data primarily for data augmentation purposes. In scenarios where sample sizes are insufficient, these studies generate additional training data to augment existing datasets (usually when the original training data is not adequate) Dai et al. (2023); Edwards et al. (2022); Møller et al. (2024); Guo et al. (2023); Ubani et al. (2023); Agrawal et al. (2023); Piedboeuf and Langlais (2023), or for generating extensive training data when high-quality human-annotated instruction tuning data is difficult to obtain Kieser et al. (2023); Taori et al. (2023); Peng et al. (2023); Xu et al. (2024). In cases where training data quality is low, many studies improve model training by providing more rationales Zhang et al. (2024); Kang et al. (2023); Li et al. (2022) or step-by-step reasoning Hsieh et al. (2023); Ho et al. (2022); Magister et al. (2023); Fu et al. (2023) to small models.
Our work distinguishes itself from existing studies by focusing on the role of familiarity in training. We explore the effectiveness of synthetic data from the perspective of familiarity, providing a theoretical foundation that substantiates the efficacy of past methods using synthetic data. Our finding offers a nuanced perspective that underscores the importance of dataset familiarity in enhancing model performance.
8 Conclusion
We systematically proved two hypotheses: LLMs are more familiar with responses generated by other LLMs when given similar inputs, and LLMs perform better on data they are familiar with. These hypotheses also explain why data annotated by LLMs yields better training outcomes than human-annotated data. The data generated by LLMs is more familiar to the trained LLMs, and training on unfamiliar data typically results in poorer outcomes. Based on this theory of familiarity, we have proposed several methods to generate data with high familiarity. Without using advanced LLMs, models can perform style transfers on ground truth to rewrite it in their own style. When using advanced LLMs for data processing, we can either use GPT-4 to read questions and generate answers in its own style, or employ our minimum change data generation method, which in some tasks, yields even better training results than direct responses from GPT-4.
9 Limitations
Perplexity can serve as a straightforward tool for measuring familiarity, but it is not very precise. It only provides relatively accurate predictions of performance decline when there are substantial differences in perplexity levels. We need a better tool to predict a model’s familiarity with a dataset and how that familiarity impacts the model’s learning.
There may be additional forms of familiarity, such as linguistic familiarity and logical familiarity. Accurately identifying and measuring the impact of these forms of familiarity presents a challenge.
Appendix A Appendix: Comparing varients of GPT-4 created training datasets
GPT-4 Answer Directly: For mathematics and coding problems, we directly present the problems from the training dataset to GPT-4 for answers. For classification tasks, we provide GPT-4 with the gold-label (which does not include human-annotated rationales) and the input, and then use its responses. We combine these answers with the original questions to create the GPT-4 Answer Training Dataset. We purposefully do not supply GPT-4 with the solutions or rationales for the math, coding, or classification problems to prevent it from merely replicating the ground truth’s problem-solving and analytical processes. If these processes were included, it could lead to a portion of GPT-4’s predictions not being generated in its own style, thus rendering our perplexity measurements—which assess how well an LLM handles predictions from other LLMs—less accurate.
Step by step transformation on GT: For classification tasks, we provide GPT-4 with both the questions and the ground truth target labels along with human-annotated rationales. For mathematical problems, we present the problems along with their answers, including the solutions process. We instruct GPT-4 to rewrite the human-annotated solutions or rationales in a step-by-step format, adding details where necessary to fill any gaps. If the existing content is already detailed enough, no additional details are added. We also instruct GPT-4 to preserve the original linguistic style as closely as possible by frequently incorporating language from the original ground truth.
Detailed Step-by-Step Transformation of Ground Truth: This data construction method is very similar to the Step by Step method. The only difference is that we specifically instruct GPT-4 to add extra details when the original ground truth lacks sufficient information.
Add details to GT: This approach involves enhancing the existing ground truth data by supplementing it with additional details to provide a richer context when necessary. We direct GPT-4 to enhance the content while maintaining the inherent language style of the original data. This method is applied selectively, only when the original information lacks sufficient clarity or depth to support effective model training and understanding.
Rewrite GT in GPT-4 style words: For this method, we present GPT-4 with the ground truth data, including human-annotated rationales and problem-solving steps. The objective is to have GPT-4 rewrite the ground truth using its own language style.
Datasets with explicit reasoning
Math Commensense Reasoning NLI
Method Training Dataset and Model Type GSM8K Math Algebra ECQA Perplexity Toekn Count
Groundtruth GSM8K, Mistral 0.434 0.162 0.594 4.72 128.882
GPT-4 Answer Directly 0.597 0.246 0.597 2.34 179.941
Rewrite GT in GPT-4 style words 0.471 0.192 0.6 4.77 156.696
Step by step 0.574 0.224 0.591 2.27 282.9
Add details to GT 0.59 0.235 0.596 3.80 272.844
Step by step transformation of GT 0.499 0.190 0.604 2.67 235.892
Detailed Step-by-Step Transformation of Ground Truth 0.506 0.196 0.602 2.59 315.028
Groundtruth GSM8K, Llama 0.364 0.141 0.565 3.27 128.882
GPT-4 Answer Directly 0.428 0.127 0.575 2.30 179.941
Rewrite GT in GPT-4 style words 0.394 0.127 0.58 3.65 156.696
Step by step 0.402 0.134 0.553 2.19 290.2
Add details to GT 0.38 0.128 0.579
Step by step transformation of GT 0.394 0.146 0.573 2.42 235.892
Detailed Step-by-Step Transformation of Ground Truth 0.396 0.114 0.553 2.37 315.028
Groundtruth Math algebra, Mistral 0.264 0.206 0.554 5.93 163.355
GPT-4 Answer Directly 0.553 0.301 0.608 2.20 328.764
Step by step 0.533 0.239 0.583 2.25 452
Add details to GT 0.445 0.230 0.57 3.148 417.35
Step by step transformation of GT 0.444 0.210 0.577 2.65 370.349
Detailed Step-by-Step Transformation of Ground Truth 0.468 0.223 0.582 2.62 484.368
Groundtruth Math algebra, Llama 0.36 0.126 0.575 5.17 163.355
GPT-4 Answer Directly 0.35 0.150 0.561 2.591 328.764
Step by step 0.344 0.144 0.593 2.55 450.1
Add details to GT 0.38 0.128 0.579 3.31 417.35
Step by step transformation of GT 0.376 0.140 0.565 2.80 370.349
Detailed Step-by-Step Transformation of Ground Truth 0.299 0.110 0.545 2.816 484.368
Groundtruth ECQA, Mistral 0.258 0.134 0.68 51.28 73.508
GPT-4 Answer Directly 0.462 0.223 0.722 6.35 176.536
Rewrite GT in GPT-4 style words 0.384 0.194 0.721 9.26 134.338
Step by step 0.481 0.203 0.71 4.66 325.2
Add details to GT 0.422 0.204 0.731 9.08 238.292
Step by step transformation of GT 0.462 0.194 0.681 6.858 218.74
Detailed Step-by-Step Transformation of Ground Truth 0.487 0.186 0.68 5.61 322.855
Groundtruth ECQA, Llama 0.132 0.079 0.631 12.16 73.508
GPT-4 Answer Directly 0.379 0.155 0.656 5.33 176.536
Rewrite GT in GPT-4 style words 0.282 0.115 0.664 6.84 134.338
Step by step 0.363 0.140 0.648 4.46 337.3
Add details to GT 0.062 0.097 0.663 8.47 238.292
Step by step transformation of GT 0.363 0.139 0.648 5.54 218.748
Detailed Step-by-Step Transformation of Ground Truth 0.135 0.106 0.66 5.08 322.855
Table 6: Appendix A: Additional experiments for Table2
Appendix B Appendix: Minimum Change Vs GPT4 answer
Datasets with explicit reasoning
Math Commensense Reasoning NLI Code
Method Training Dataset and Model Type GSM8K Math Algebra ECQA ESNLI HumanEval MBPP Perplexity Toekn Count
Groundtruth GSM8K, Mistral 0.434 0.162 0.594 0.597 4.53 128.88
GPT-4 Answer Directly 0.597 0.246 0.597 0.587 2.34 179.941
Mistral Minimum Change Data 0.562 0.234 0.597 0.558 1.67 177.156
llama minimum change data 0.547 0.242 0.621 3.19 166.152
Groundtruth GSM8K, Llama 0.371 0.143 0.566 0.559 3.27 130.856
GPT-4 Answer Directly 0.428 0.127 0.575 0.568 2.30 183.852
Llama Minimum Change Data 0.433 0.140 0.602 0.602 2.22 166.15
Mistral Minimum Change Data 0.402 0.141 0.576 2.249 177.156
Groundtruth Math algebra, Mistral 0.264 0.206 0.554 0.574 5.83 163.355
GPT-4 Answer Directly 0.553 0.301 0.608 0.555 2.19 328.764
Mistral Minimum Change Data 0.536 0.313 0.622 0.558 1.85 297.609
llama minimum change data 0.546 0.296 0.63 2.47 269.688
Groundtruth Math algebra, Llama 0.36 0.126 0.575 0.552 5.14 162.417
GPT-4 Answer Directly 0.35 0.150 0.561 0.564 2.59 326.573
Llama Minimum Change Data 0.412 0.166 0.602 0.554 2.268 269.68
Mistral Minimum Change Data 0.32 0.160 0.568 2.53 297.609
Groundtruth ECQA, Mistral 0.258 0.1348 0.68 0.544 51.29 73.508
GPT-4 Answer Directly 0.462 0.223 0.722 0.589 6.35 176.53
Mistral Minimum Change Data 0.433 0.192 0.699 0.542 3.798 112.591
LLama Minimum Change Data 0.483 0.212 0.709 4.89 128.695
Groundtruth ECQA, Llama 0.132 0.079 0.631 0.431 11.95 76.118
GPT-4 Answer Directly 0.379 0.155 0.656 0.479 5.33 183.182
Llama Minimum Change Data 0.392 0.128 0.649 0.522 3.619 128.695
Mistral Minimum Change Data 0.28 0.126 0.647 5.58 112.591
Groundtruth ESNLI, Mistral 0.417 0.172 0.594 0.814 378.7 29.925
GPT-4 Answer Directly 0.464 0.214 0.592 0.804 6.54 137.967
Mistral Minimum Change Data 0.453 0.206 0.602 0.831 3.38 82.367
Llama Minimum Change Data 0.432 0.203 0.63 0.836 4.15 97.078
Groundtruth ESNLI, Llama 0.332 0.125 0.582 0.751 1201.70 29.963
GPT-4 Answer Directly 0.386 0.123 0.564 0.776 5.28 143.739
Llama Minimum Change Data 0.357 0.128 0.578 0.796 3.74 97.078
Mistral Minimum Change Data 0.346 0.130 0.546 0.811 4.34 82.367
Groundtruth HumanEval, Mistral 0.362 0.191 0.583 0.531 0.323 5.69 71.60
GPT-4 Answer Directly 0.313 0.162 0.581 0.442 0.365 4.71 107.9
Mistral Minimum Change Data 0.4055 0.202 0.5915 0.541 0.408 2.21 109.073
Llama Minimum Change Data 0.4115 0.179 0.5835 0.5505 0.426 2.451 122.0
Groundtruth HumanEval, Llama 0.0705 0.083 0.528 0.339 0.146 5.75 70.8
GPT-4 Answer Directly 0.1245 0.105 0.553 0.473 0.158 4.33 107.14
Llama Minimum Change Data 0.2445 0.389 0.5705 0.5175 0.213 2.21 122.0
Mistral Minimum Change Data 0.131 0.103 0.528 0.4235 0.183 2.75 109.07
Groundtruth MBPP, Mistral 0.392 0.175 0.527 0.481 0.276 7.78 71.60
GPT-4 Answer Directly 0.399 0.186 0.568 0.553 0.354 3.419 786
Mistral Minimum Change Data 0.405 0.179 0.601 0.48 0.354 2.817 63.86
Llama Minimum Change Data 0.332 0.179 0.56 0.511 0.364 4.55 54.96
Groundtruth MBPP, Llama 0.328 0.131 0.556 0.537 0.2 4.74 73.90
GPT-4 Answer Directly 0.351 0.139 0.574 0.553 0.202 3.33 69.78
Llama Minimum Change Data 0.34 0.130 0.569 0.539 0.204 3.44 54.96
Mistral Minimum Change Data 0.333 0.123 0.563 0.544 0.218 3.45 63.86
Table 7: Appendix B: More results for Table 5. Comparing the experimental results of GPT4, groundtruth and minimum change
Appendix C Appendix: Minimum Change Prompt and Method
Figure 3: Appendix C: Minimum Change Data Correction Examples
Its construction process is as follows: For a given question, we first ask the target LLM, such as Llama2 or Mistral, to produce a response. If the response is correct, we include it in the synthetic dataset. If not, we instruct GPT-4 to revise it, making minimal changes to the original answer to correct it  [^2] Please note that achieving such a process requires an advanced LLM, as the LLM needs to fully understand the solution and follows the “minimal change” principal. Note that this construction process ensures that the target LLM is “familiar” with the response as the response is largely generated from itself. The “minimal change” principal, on the other hand, minimizes the influence from the GPT-4.
Specifically, minimal change is implemented via the following steps:
  • Initial Prediction the model generates an initial prediction.
  • Minimal Modification uses GPT-4 to adjusts the initial prediction with the fewest possible changes to correct errors, deliberately preserving the original distribution by keeping as many words unchanged as possible. An illustration of such a modification is provided in Figure 2. We have detailed the prompt used to guide GPT-4 in the ’Minimum Change Prompt’ section of the Appendix.
  • Training with Modified Predictions the minimum changed predictions are subsequently used as target labels for model training.
Figure 4: Appendix C: Minimum Change Prompt Example
This method involves minimal intervention by GPT-4 in the initial predictions to ensure that the original language style is largely retained. To effectively guide GPT-4 in producing these minimally changed training datasets, we used a specific prompt, outlined in the ’Minimum Change Prompt’ section of the Appendix, which includes two to three examples of minimal changes along with explanations for each modification.
Appendix D Appendix: LLaMA groundtruth style transfer failure example
The LLaMA-13b-chat model often fails to follow our in-context examples, leading to incorrect outputs when it is tasked with modifying ground truth data.
Figure 5: Appendix D: LLaMA groundtruth style transfer failure example